Consider a 60.0 kg box that is pushed up an 8.0 m long ramp with a force of 100.0 N to a height of 1.0 m in 5 seconds. What is the efficiency % of the ramp?
Answer #1

The\; formula \; of \; efficiency\; (\gamma_{percent} )\; =\;( W_{out}/W_{in})\times 100

W_{out}=Work \; done\; by \; output\; force

W_{in}=Work \; done\; by \; input\; force

W_{out}=mgh

W_{out}=60\times 9.8\times 1

\mathbf{W_{out}=588J}

\mathbf{Now\; we \; need \; to\; calculate W_{in}}

W_{in}}=Fd

Where \; F \; is \; the \; applied \; force\; on \; the \; block

And \; d=\; displacement

W_{in}=Fd

W_{in}=100N\times 8m

W_{in}=800\; \; Nm

\mathbf{Now \; use\; the\; formula\; of\; efficiency(\gamma_{percentage} )=(W_{out}/W_{in}})\times 100

                                                             efficiency(\gamma_{percentage} )=(588J/800J)\times 100

                                                               efficiency(\gamma_{percentage} )=(.735)\times 100

                                                                \mathbf{efficiency(\gamma_{percentage} )=(73.5)\; \; Percentage}             (Ans)